OpenStudy (anonymous):

what is the simplified form of sqrt{162}

OpenStudy (anonymous):

162 = 2 * 81 sqrt 162 = sqrt(2*81) = 9 sqrt2

OpenStudy (lgbasallote):

factor into prime numbers \(\large \sqrt{(2)(3)(3)(3)(3)}\) take out of the square root all that appeared twice